OpenStudy (rational):

show that \(m^2+n+2\) and \(n^2 +4m\) are not perfect squares simultaneously

OpenStudy (amistre64):

limited to integers? whole numbers?

OpenStudy (rational):

sry yes limited to whole numbers

OpenStudy (rational):

*positive integers

OpenStudy (fibonaccichick666):

oooh! hey I can actually do this one!

OpenStudy (rational):

i bet you can :)

OpenStudy (fibonaccichick666):

Number theory proofs are not my strong suit, so I get excited when I can as it is not a usual occurrence.

OpenStudy (fibonaccichick666):

So, I would approach this by first defining what it would mean if each respectively was a perfect square

OpenStudy (rational):

here is a hint given book if \(k^2\) is a perfect square (it is!) then the next perfect square occurs after \(2k+1\) integers that means for \(m^2+n+2\) to be a perfect square, we must have \(n+2 \ge 2m+1\)

OpenStudy (fibonaccichick666):

I think you can pull a wlog m>n

OpenStudy (dan815):

true

OpenStudy (fibonaccichick666):

as for the hint, do you see where they got that from?

OpenStudy (fibonaccichick666):

not you dan haha

OpenStudy (rational):

not quite.. im still unsure of this prblem actually

OpenStudy (fibonaccichick666):

ok so let's look at it from the beginning. Do you agree that if \(m^2+n+2=k^2 \) is a perfect square then there exists some k>2 for it?

OpenStudy (fibonaccichick666):

we can actually make that sharper I think

OpenStudy (amistre64):

a^2 - b^2 = (a+b)(a-b) a^2 = m^2+n+2 -b^2 = -(n^2+4m) m^2 +n +2 -n^2 -4m (m^2-4m) - (n^2 -n) +2 (m^2-4m+4) - (n^2 -n+1/4) +2-4+1/4 (m-2)^2 - (n-1/2)^2 -7/4 = (m-2+n-1/2)(m-2-n+1/2) -7/4 not sure where im going with this thought .... just thinking tho

OpenStudy (fibonaccichick666):

ugh, but we need to prove the perfect squares layout first in a lemma

OpenStudy (dan815):